OpenStudy (firejay5):

What are the 4 corners of the dotted box of this hyperbola? Give the coordinate points. Already given: (5, - 8) & (5, 12) <---- part of graphing hyperbola

OpenStudy (firejay5):

\[\frac{ (y-2)^2 }{ 100 } - \frac{ (x-5)^2 }{ 4 } = 1\]

OpenStudy (firejay5):

The Asymptote of the hyperbola is: y = 2 +/- 5 (x - 5)

OpenStudy (firejay5):

@cwrw238 @ganeshie8 @satellite73

OpenStudy (paxpolaris):

\[(5\pm2,2\pm10)\]

OpenStudy (firejay5):

okay wait what??? Center is at (5, 2)

OpenStudy (firejay5):

I need like 2 up and down from center for example

OpenStudy (paxpolaris):